OpenStudy (anonymous):

Suppose that a pair of dice is tossed. One die has 8 sides, the other has 6 sides. What is the expected value of the sum of the two dice?

OpenStudy (kropot72):

The sample space has 48 possible outcomes. You need to find the 48 values of sums and add them. Then divide the total by 48. What you are doing is finding the average value of the 48 values of sums.

OpenStudy (anonymous):

sum = 2: 1/1, total = 1 sum = 3: 1/2, 2/1, total = 2 sum = 4: 1/3, 2/2, 3/1, total = 3 sum = 5: 1/4, 2/3, 3/2, 4/1, total = 4 sum = 6: 1/5, 2/4, 3/3, 4/2, 5/1, total = 5 sum = 7: 1/6, 2/5, 3/4, 4/3, 5/2, 6/1, total = 6 sum = 8: 1/7, 2/6, 3/5, 4/4, 5/3, 6/2, total = 6 sum = 9: 1/8, 2/7, 3/6, 4/5, 5/4, 6/3, total = 6 sum = 10: 2/8, 3/7, 4/6, 5/5, 6/4, total = 5 sum = 11: 3/8, 4/7, 5/6, 6/5, total = 4 sum = 12: 4/8, 5/7, 6/6, total = 3 sum = 13: 5/8, 6/7, total = 2 sum = 14: 6/8, total = 1 The average value of the 48 values is 8. Thank you!

OpenStudy (kropot72):

You're welcome :)